# Review, Refine, Repeat: Understanding Iterative Decoding of AI Agents with Dynamic Evaluation and Selection

### v1

@@ B.2 Sketch2code
For Sketch2code (Li et al., 2024b), we provide a detailed comparison of our approach against SoTA baselines on several evaluation criterion and metrics. We used the hyperparameter setting of temperature = 0.6, max tokens = 4096, top p = 1.0, frequency/repetition penalty = 0.0, and presence penalty = 0.0 for all our results. For the metrics, we consider metrics centring 1.Layout Similarity, 2. Visual IoU, 3. Text IoU with reference HTML following (Li et al., 2024b). These metrics offer a comprehensive and reliable assessment of HTML generation quality, demonstrating a strong correlation (9̃0%) with human satisfaction, as shown in (Li et al., 2024b) (further details in Appendix). Hence, we use Layout similarity as a verifier along with LLM-as-judge (Li et al., 2024b) to guide the generations for both BON (Beirami et al., 2024) and IAD. We report comparison with baseline single-turn approaches including SotA models GPT-4o, Claude-3, InternVL2, Gemini-1.5-Flash, CoT and variants along with multi-response generation approaches including BON, Sk2code and IAD (Ours). Single turn approaches even from SoTA models fail to match the layout structure, position of blocks, textual content, size of the blocks etc in the given user-sketch, causing a mismatch w.r.t to the reference layout as can be clearly seen in Figure and achieves a low score in-terms of all the three metrics in-comparison with multi-response generation approaches even with N=2. Best-of-N sampling (BON) with a weaker model Gemini-1.5-Flash improves over single-turn approaches and , with $N=4$ generations, it outperforms SoTA models with single-turn responses by a margin of 15-18%, by correctly identifying the block position, title block, overall layout structures etc. We see monotonic improvement in performance over the number of responses as the layout score improves from $20.41$ to $25.7$ with 6 responses. However, BON struggles in incorporating fine-grained details about layout structure and makes some-times makes repetitive mistakes in the position of block in all the $N$ generation for the prompts (as shown in Fig). Our proposed approach IAD, mitigates this gap by iteratively improving the responses and as shown in Table 1, it achieves a major improvement of $\leavevmode\nobreak\ 3-5\%$ (absolute improvement) from BON as well as single-turn SoTA Claude with just 2 iterations (eq : N=2) even with simpler model Gemini-1.5-Flash. At each iteration, we pass the best and worst HTML as a context along with instruction, for generating the next iteration. We observe IAD is able to learn fine-grained layout components, image semantics over iterations with the context of the Best and Worst HTML. We see that with increased iterations, performance of IAD improves reaching to a very high layout score of 26.75, outperforming all baselines with same generations. We also report the Image and Text IoU scores while optimizing with the layout-score, to check for reward-overoptimization of the metric.

### v4

@@ B.2 Sketch2code
For Sketch2code (Li et al., 2024b), we provide a detailed comparison of our approach against SoTA baselines on several evaluation criterion and metrics. We used the hyperparameter setting of temperature = 0.6, max tokens = 4096, top p = 1.0, frequency/repetition penalty = 0.0, and presence penalty = 0.0 for all our results. For the metrics, we consider metrics centring 1.Layout Similarity, 2. Visual IoU, 3. Text IoU with reference HTML following (Li et al., 2024b). These metrics offer a comprehensive and reliable assessment of HTML generation quality, demonstrating a strong correlation (9̃0%) with human satisfaction, as shown in (Li et al., 2024b) (further details in Appendix). Hence, we use Layout similarity as a verifier along with LLM-as-judge (Li et al., 2024b) to guide the generations for both BON (Beirami et al., 2024) and IAD. We report a comparison with baseline single-turn approaches, including SotA models GPT-4o, Claude-3, InternVL2, Gemini-1.5-Flash, CoT, and variants, along with multi-response generation approaches, including BON, Sk2code, and IAD (Ours). Single turn approaches even from SoTA models fail to match the layout structure, position of blocks, textual content, size of the blocks etc in the given user-sketch, causing a mismatch w.r.t to the reference layout as can be clearly seen in Figure and achieves a low score in-terms of all the three metrics in-comparison with multi-response generation approaches, even with N=2. Best-of-N sampling (BON) with a weaker model Gemini-1.5-Flash improves over single-turn approaches and, with $N=4$ generations, it outperforms SoTA models with single-turn responses by a margin of 15-18%, by correctly identifying the block position, title block, overall layout structure, etc. We see monotonic improvement in performance over the number of responses as the layout score improves from $20.41$ to $25.7$ with 6 responses. However, BON struggles in incorporating fine-grained details about layout structure and makes some-times makes repetitive mistakes in the position of block in all the $N$ generation for the prompts (as shown in Fig). Our proposed approach IAD, mitigates this gap by iteratively improving the responses and as shown in Table 1, it achieves a major improvement of $\leavevmode\nobreak\ 3-5\%$ (absolute improvement) from BON as well as single-turn SoTA Claude with just 2 iterations (eq : N=2) even with simpler model Gemini-1.5-Flash. At each iteration, we pass the best and worst HTML as a context along with instructions for generating the next iteration. We observe IAD is able to learn fine-grained layout components, image semantics over iterations with the context of the Best and Worst HTML. We see that with increased iterations, the performance of IAD improves, reaching a very high layout score of 26.75, outperforming all baselines with the same generations. We also report the Image and Text IoU scores while optimizing with the layout-score, to check for reward-overoptimization of the metric.